THE AMOUNT A IN AN ACCOUNT AFTER t YEARS FROM AN INITIAL PRINCIPLE P INVESTED AT AN ANNUAL RATE r COMPOUNDED CONTINUOUSLY IS GIVEN BY A =Pe^rt WHERE r IS EXPRESSED AS A DECIMAL. SOLVE THIS FORMULA FOR t IN TERMS OF A,P,AND r.
Answer by stanbon(48568)

A =Pe^(rt)
A/P = e^(rt)
ln(A/P)=rt
t=(1/r)[ln(A/P)]
------------
Cheers,
Stan H.